OpenStudy (anonymous):

whats the difference. (6/x^2-4)-(4/x^2-16)

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

\[\frac{ 6 }{ x^2-4 }-\frac{ 4 }{x^2-16 }=\frac{ 6\left( x^2-16 \right)-4\left( x^2-4 \right) }{ \left( x^2-4 \right)\left( x^2-16 \right) }\] \[=\frac{ 6x^2-96-4 x^2+16 }{ \left( x^2-4 \right)\left( x^2-16 \right) }=?\]